Project Cost= 10,431.70 Life of the plant= 12 year Salvage Value= 521.59 EBIT= 1877 Tax Rate= 35% WACC= 5.73% NPR? IRR? Discounted payback period ?
Answer #1

Depreciation per year = (10,431.70 - 521.59) / 12
= 825.84

EBIT = 1877
Less: Tax 35% = 656.94
Net Income = 1220.05
Add: Depreciation = 825.84
Annual cash inflow = 2045.89

Calculation of NPV and IRR.

Here IRR is calcualted using the excel formula of IRR.

Discounted payback period:
Total present value of cash inflows upto 6 years = $10,145.98
Balance of Investment will be recovered in Year 7 i.e. $10,431.70 - $10,145.98 = $285.72
Discounted payback period = 6 years + $285.72 / $1,385.07
= 6 + 0.21 = 6.21 years